x?y + Oxy + 9xy + (16 + 2)y = 0, y(1) = 0, yle7/9) = 0

(a) Find the eigenvalues.

as a symbolic function of n

(b) Find the eigenfunctions. Take the arbitrary constant (either c1 or c2) from the general solution to be 1.

as a symbolic function of x,n
Answer #1

Given equation is x²y+any+ (16+) y =0 Let x-et at is the new independent variable. then xy = D(D-1) Y D=d JE axy 30, © becoO= 4+ g = G = -2 Now, y(1)=0 gives y (eda)=0 (4+5 gives 0 = ce 9 (+4-FT/ the (4-5-x) Na 7 C X7o, 64+50W 645) los te 4 = 0 po4t case 3: then the solection of ② is y = 4 sin (JFX t) + Co Cos (Stit) é y@ = G Sin (Nar Inces) + (n_C06 (J73 In ()} & 1 Now